From GlueXWiki
Jump to: navigation, search

FDC Weekly Meeting

Date: June 18, 2009

Participants: Daniel, Simon, Brian, Micah, Rich, Bill, Mark S., Beni, Fernando, Roger, Chuck

Next Meeting: June 25, 2009 @ 1:30 p.m.

General Construction Updates

 - The fiber-reinforced gusset pieces have arrived from the shop. Each ring
   will be made from 8 arc pieces, with one ring on the upstream package
   face and one ring on the downstream package face.
 - The parts for the assembly jig on which the detector stack will be
   constructed are now being made in the machine shop. Their delivery is
   imminent. Once we have the pieces in hand, we will do some mechanical
   checks to see how everything fits together.
 - Bill will take on the preparation for a stack assembly document. A
   draft will be prepared before assembly of the full-scale prototype and
   then fleshed out during the process with relevant photographs taken.
 - Cable preparations:
    * The LV cables have arrived and will be tested in the near future.
    * The parts n' pieces for the HV cables are now all in hand and
      construction and testing will begin over the next several weeks.
    * The signal cables and connectors are in hand and Mark and Brian will
      begin construction as they have time.
    * We discussed how much of the outer jacket and braid will be stripped
      off the signal cables. The initial thought was to pull it off all the
      way to the upstream face of the magnet. We need to get some feedback
      from the BCAL folks on this. Fernando will post some information on
      this question on the wiki.
 - The o-ring material order has arrived. The material is a mixture of
   nitrile and EPDM (Ethylene Propylene). Its hardness is about
   10 durometers.
 - We are now having short daily meeting to review the work on the
   full-scale prototype construction and to ensure that any snags or
   problems that come up are immediately addressed.
 - We discussed the time frame for preparing the upstream and downstream
   gas windows and their associated laminate frames. Bill thought that the
   laminations could be laid up next week and set aside for when we are
   ready to attach the aluminized mylar windows.
 - Mark S. will prepare the order for the cables that will be used to
   attach to the cathodes for the external ground. He will also order the
   sealing epoxy that is used to cover the conducting epoxy. This is a
   humidity protection plan recommended by Fernando.
 - Eugene is working to identify clean room space for FDC construction.
   He has talked recently with Dennis Skopik who will work to define the
   assigned space.

Wire Frame Update

 - Armen has completed the soldering work on the four wire frames. All four
   frames were placed back in the Lucite box under nitrogen purge.
 - Rich has been performing HV tests on the frames one by one. Three frames
   have been tested with one to finish. The test amounts to applying 1kV
   to each bus and studying the current draw. A problem has been found with
   one bus on frame P3. Fernando and Rich will come back to this board and
   try to isolate the cause. More to come. Note that this electrical QA
   on the finished frames will be a standard part of QA on all wire frames.
 - We need to have a dedicated meeting to design what the schedule for
   work on the Phase 3 wire winding will be. The highest priority for
   work scheduling will be doing work on upgrades of the wire winding
   facility. This includes a better strongback design, the design and
   construction of precision combs, modification to the wire winding
   table, and a list of other items with the system. Stay tuned, more
   to come.
 - We are pushing hard to begin wire frame construction early next year and
   begin wire winding before the end of 2010. Hopefully this can all procede.
   We are now working to complete the simulation studies with Simon to
   determine if we want to offset neighboring wire layers. More to come in
   the upcoming months.
 - The wire frame construction document is located at:
   /u/group/halld/Engineering/PRELIMINARY DOCUMENTS. He will include
   the information on the wire electroplating process as well.
 - Bill is now designing the jig for the wire electroplating operation
   and it will be ready for submission to the shop at some point in the
   future. We will make a decision shortly if/when we will electroplate
   the wires for the full-scale prototype.
 - Brian received the SEM scans for our wire from the W&M folks. He will
   generate a final QA report with photos and test results for all wires
   in the next several weeks. 

Cathode Update

 - Progress updates:
     * The laminations for the 8 cathode frames have been completed. Work 
       to seal the exposed Rohacell edges and drill and seal the gas ports
       is now underway. This is a relatively slow process that we can speed
       up with a bit of cleverness.
     * Cutting of the edges of the cathode boards is now nearly finished.
       Brian and Mark S. should be done by tomorrow with the full set.
     * We are working to finalize the procedure for laying up the 3 piece
       cathode panels. It appears we have a complete plan in place and
       a final mockup has been made with dummy boards. 
     * A cathode panel lay up using a non-final procedure was completed
       and the quality is not superior, but it should be useable. The
       free transfer ring has been attached and the board is ready for
     * Brian and Mark S. expect to lay up the next cathode panel tomorrow
       and should be able to complete 2 or 3 next week.
     * Eugene has retensioned the 2 micron board that was made a month ago.
       He set the average tension to ~500 N/m and saw a 2% tension change
       after 1+ day. Measuring 5 days later found no further tension
       changes. This board should have a frame attached this week and
       removed from the tensioning facility.
     * Micah and Mark S. have scanned the surface flatness of the 5 micron
       Cu board and Micah showed us some preliminary results that were
       encouraging. He wants to scan the granite table and subtract this
       data to give the final flatness measurements for this board. The
       2 micron Cu board will be scanned next week. Note that this work
       occurred only after Bill spent significant time modifying the system
       to ensure the cathode boards were clamped flat to the granite table.
 - Brian will construct a bag that we can keep all completed cathodes in
   so that we can keep them in a nitrogen purge to prevent oxidation. The
   bag will allow the completed cathodes to be clamped so that the frames
   remain flat.
 - A portion of the rigid-flex assembly has arrived (50 each of the cathode
   boards and ground boards). The problem with the locating holes being to
   small (mentioned above for the wire frame PCBS) is present with these
   boards too. Roger will see to it that the holes are enlarged and the
   boards are passed on to Armen for soldering work early next week. Fernando
   has already alerted Armen so that he is ready to receive the boards. The
   remainder of the order will be delivered in about 2 weeks.
 - Bill has been working on the cathode construction document. He has
   circulated a draft for comment. The file is located at
   /u/group/halld/Engineering/PRELIMINARY DOCUMENTS.
 - We are working to finalize the PR for the 2 micron cathode board
   material. Procurement will be negotiating with Nippon Steel Chemical
   Company to try to convince them to complete this order. Roger is also
   in discussion with the company that makes the 5 micron Cu boards about
   making the 2 micron boards. Stayed tuned for new developments.
 - Roger needs to prepare a document for QA/stuffing/cleaning for the
   cathode boards and a similar document for the cathode daughter
   boards and ground boards.

Cooling System Tests

 - Fernando is in the process of writing up the results from his cooling 
   system studies.  Stay tuned for the GlueX note. He indicated that he
   is back to work on this document and it should be done shortly.

Cosmic Ray Test Stand

 - Beni is continuing work to setup the cosmic ray telescope and test
   system in EEL 126. He would like to have it all up and functioning
   and understood before the end of June. Progress is being made connecting
   up all detectors and getting CODA working.
 - We still have some mysterious behavior with the small-scale prototype
   response with wiggles in the X vs. Y distributions that need to be
   understood. We decided to have a separate meeting to discuss what work
   we would like to do on the small-scale prototype.
 - Beni has started a bench test plan for the full-scale prototype and
   we discussed his preliminary document briefly. He will circulate it
   for feedback and we will continue to flesh it out. We want to make sure
   that we do some checks of gas leakage with the prototype. We agreed that
   the best procedure would be to use the P5 gas mixture (non-flammable).
 - Bill has designed a system to rotate the prototype from vertical to 
   horizontal. The drawings have been submitted to the machine shop and the
   pieces should be delivered in about 3 weeks.
 - Safety paperwork for the cosmic ray test stand is being prepared now
   and will be submitted to Bert shortly.

Work List

 - The FDC short-term work list has been posted on the FDC web site.  
   This is continually being updated and DSC welcomes any feedback or 
   comments from the group.

Minutes prepared by Daniel. Send any comments or corrections along.